OpenStudy (anonymous):

differentiate y=2x^4

OpenStudy (lgbasallote):

\[\large x^n = nx^{n-1}\]

OpenStudy (apoorvk):

same as the prev. one. \[\large \frac {d(x^n)}{dx} = nx^{n-1}\]

OpenStudy (anonymous):

is the answer 8x^3

OpenStudy (apoorvk):

yes, right-o!

OpenStudy (anonymous):